what is the answer of x(x+3)+2(6x-5)=10?

Answer by checkley71(8403)

X(X+3)=2(6X=5)=10
X^2+3X+12X-10=10
X^2+15X-10-10=0
X^2+15X-20=0
USING THE QUADRATIC EQUATION x = (-b +- sqrt( b^2-4*a*c ))/(2*a)
X=(-15+-SQRT[15^2-4*1*-20])/281
X=(-15+-SQRT[225+80])/2
X=(-15+-SQRT305)/2
X=(-15+-17.46)/2
X=(-15+17.46)/2
X=-2.46/2
X=-1.23 ANSWER.
X=(-15-17.46)/2
X=-32.46/2
X=-16.23 ANSWER.